OpenStudy (anonymous):

can you tell me if B is right? Which line has an undefined slope? A. y = 0 B. y = 2 C. x = y D. x = 2

OpenStudy (anonymous):

I mean A not B

OpenStudy (dls):

Yes

OpenStudy (dls):

\[\LARGE \tan \theta=\frac{y}{x}\]

OpenStudy (hba):

None of these...a slope that is undefined is a line that is vertical. It's easy to remember because anything that you could possibly walk on has a slope. You can't walk up a wall (vertical), so the slope is undefined (or you can say it has no slope). Keep in my mind that a horizontal line (like the floor) has a slope; the slope is zero. Undefined = No Slope!!!!! Slope = 0 has a slope, the slope is zero. -Thanks to : Manders.

OpenStudy (dls):

\[\LARGE m=\frac{0}{x}\]

OpenStudy (anonymous):

okay thanks

OpenStudy (dls):

Its still not undefined but its 0.. if x was 0 we cud say undefined